the unit's digit of a number is 3 more then the ten's digit of it. On reversing the order of the digits, the numbers obtained is 2 more then the double of the original number. what is the sum digits of the original number?
Answer by lwsshak3(11628)

the unit's digit of a number is 3 more then the ten's digit of it. On reversing the order of the digits, the numbers obtained is 2 more then the double of the original number. what is the sum digits of the original number?
***
let u=units digit
let t=tens digit
..
u=t+3
(10u+t)=2(10t+u)+2
10t+30+t=20t+2u+2
11t+30=20t+2t+6+2
11t=22
t=2
u=t+3=5
what is the sum digits of the original number?
units digit=5
tens digit=2